The area of a triangle is 34 square meters. Find the base and height of the latter exceeds the former by 67 meters.
Please help..thanks..

Answer by scott8148(6628)

let x=base, x+67=height ... (1/2)(x)(x+67)=34 ... x(x+67)=68 ... x^2+67x-68=0

(x+68)(x-1)=0 ... x=-68 and x=1 ... negative value not realistic

base=1, height=68